As part of maintenance, the StanCam and FIES pick-off mirrors where cleaned with alcohol. It was noted that the FIES mirror shows wear, and a large stain. One of the spare FIES pick-off mirrors was re-aluminized by ING together with our main mirror. The resulting reflectivity is not higher than that of new mirrors and we bought new mirrors from Edmund and measured a reflectivity of 93%,93%,92%,81% in B, V, R and z, respectively. The plan is to replace the mirror in the near future.

The upgrade of the original fiber bundle has proceeded and was refurbished at ESO by Gerardo Avila. The University of Aarhus provided new mechanical parts for the fiber couplings. The ``new'' bundle was received at NOT, and has been installed from the hole under the telescope to the FIES spectrograph. All 4 fibers transmitted healthy amounts of light after the installation. A new clamp to hold the new fiber bundle at the spectrograph entrance was made to account for the fact that the fibers are wrongly fixed and rotated with respect to the optical axis. The testing and subsequent commissioning of the new bundle is planned in the coming months.

To make any alignment of the fiber easier after any misalignment an XY-slide has been purchased and prepared for installation. This is to ensure proper alignment between the adapter-based calibration fibers and the main astronomical fibers. The slide will be mounted in the adapter plate. The adapter plate itself will be move slightly to allow the position of all fibers to be within the FOV of StanCam.

A continuous check of the power-on status of the top-calibration unit electronics was implemented. If the power is found to be off, the staff will be automatically notified to initialize the box.

The FIES spectrograph itself was refocused, showing a change in focus position from 196 to 198 with no significant change in PSF. The optimal telescope focus for each fiber still needs to be tested by checking count rates at FIES detector as a function of telescope focus.

Thomas Augusteijn 2010-02-09